Consider the following vectors W = {1+ x, 1-2x, 3+x}

a) Test or refute if the set W is linearly independent.

b) Build a base B for polynomials of degree less than or equal to 1 (p(x) = a+bx),

c) Write the matrix associated with base B of part b).

d) Vector representation of the polynomial f(x)=x in the base B of part b).
Answer #1

Wp 3 (148), (-24), (6+2) 4 Jet 2(142) + B (1-2x) + r (3+x7 zo (943438) + X(X-26+Y)=0 a 2+3+3r=0 (4-23+v)=0 Solving from and x© Matin representation Associated with Basis 21,09 110 fo.11 __ =9.(1? + 1.36 . Mahia becomes O, TO @ vectos representation o